# MInCo: Mitigating Information Conflicts in Distracted Visual Model-based Reinforcement Learning

### v1

@@ IV-A Information Conflicts in Visual MBRL
$\mathbb{E}$ $[\text{KL}(p(s_{t}|o_{t},s_{t-1},a_{t-1}\parallel q(s_{t}|s_{t-1},a_{t-1}))]$
@@ IV-E Policy Learning
$\mathcal{J}_{\mathrm{MInCo}}\doteq\mathbb{E}_{p}\big{(}{\sum_{t}\Big{(}\mathcal{J}_{\mathrm{SimSiam}}^{t}+\mathcal{J}_{\mathrm{R}}^{t}+\mathcal{J}_{\mathrm{D_{tv}}}^{t}+\mathcal{J}_{\mathrm{c-inv}}^{t})}).$ (15)
@@ Appendix A Proofs for Proposition 1 and 2
The complete derivation of Proposition 2 is as follows: Proof.
$I(s_{t},o_{t}|s_{t-1},a_{t-1})$
$\equiv\mathbb{E}[\ln\frac{p(s_{t}|o_{t},s_{t-1},a_{t-1})}{p(s_{t}|s_{t-1},a_{t-1})}]$
$=\mathbb{E}[\ln\frac{p(s_{t}|o_{t},s_{t-1},a_{t-1})q(s_{t}|s_{t-1},a_{t-1})}{q(s_{t}|s_{t-1},a_{t-1})p(s_{t}|s_{t-1},a_{t-1})}]$
$=\mathbb{E}[\ln\frac{p(s_{t}|o_{t},s_{t-1},a_{t-1})}{q(s_{t}|s_{t-1},a_{t-1})}]+\mathbb{E}[\ln\frac{q(s_{t}|s_{t-1},a_{t-1})}{p(s_{t}|s_{t-1},a_{t-1})}]$
$=\mathbb{E}[\ln\frac{p(s_{t}|o_{t},s_{t-1},a_{t-1})}{q(s_{t}|s_{t-1},a_{t-1})}]$
$-\mathbb{E}[\text{KL}(p(s_{t}|s_{t-1},a_{t-1})\parallel q(s_{t}|s_{t-1},a_{t-1}))]$
$=\mathbb{E}[\ln\frac{p(s_{t}|o_{t},s_{t-1},a_{t-1})}{q(s_{t}|s_{t-1},a_{t-1})}]$
$-\text{KL}(p(s_{t}|s_{t-1},a_{t-1})\parallel q(s_{t}|s_{t-1},a_{t-1}))$
$\leq\mathbb{E}[\text{KL}(p(s_{t}|o_{t},s_{t-1},a_{t-1})\parallel q(s_{t}|s_{t-1},a_{t-1}))].$

### v2

@@ IV-A Information Conflicts in Visual MBRL
$\mathbb{E}$ $[\text{KL}(p(s_{t}|o_{t},s_{t-1},a_{t-1})\parallel q(s_{t}|s_{t-1},a_{t-1}))]$
@@ IV-E Policy Learning
$\mathcal{J}_{\mathrm{MInCo}}\doteq\mathbb{E}_{p}\big{(}{\sum_{t}\big{(}\mathcal{J}_{\mathrm{SimSiam}}^{t}+\mathcal{J}_{\mathrm{R}}^{t}+\mathcal{J}_{\mathrm{D_{tv}}}^{t}+\mathcal{J}_{\mathrm{c-inv}}^{t}\big{)}}\big{)}.$ (15)
@@ Appendix A Proofs for Proposition 1 and 2
The complete derivation of Proposition 2 is as follows:
$I$ $(s_{t},o_{t}\mid s_{t-1},a_{t-1})$
$=\mathbb{E}_{p(s_{t},o_{t}\mid s_{t-1},a_{t-1})}\left[\ln\frac{p(s_{t}\mid o_{t},s_{t-1},a_{t-1})}{p(s_{t}\mid s_{t-1},a_{t-1})}\right]$
$=\mathbb{E}\left[\ln\frac{p(s_{t}\mid o_{t},s_{t-1},a_{t-1})}{q(s_{t}\mid s_{t-1},a_{t-1})}-\ln\frac{p(s_{t}\mid s_{t-1},a_{t-1})}{q(s_{t}\mid s_{t-1},a_{t-1})}\right]$
$=\mathbb{E}_{p(o_{t}\mid s_{t-1},a_{t-1})}\left[\mathrm{KL}\left(p(s_{t}\mid o_{t},s_{t-1},a_{t-1})\,\|\,q(s_{t}\mid s_{t-1},a_{t-1})\right)\right]$
$\quad-\mathrm{KL}\left(p(s_{t}\mid s_{t-1},a_{t-1})\,\|\,q(s_{t}\mid s_{t-1},a_{t-1})\right)$
$\leq\mathbb{E}_{p(o_{t}\mid s_{t-1},a_{t-1})}\mathrm{KL}\left(p(s_{t}\mid o_{t},s_{t-1},a_{t-1})\,\|\,q(s_{t}\mid s_{t-1},a_{t-1})\right)$